OpenStudy (anonymous):

The sum of twice a number and 12 less than the number is the same as the as the difference between -28 and the number. What is the number?

OpenStudy (whpalmer4):

twice a number: 2x 12 less than the number: x-12 sum of twice a number and 12 less than the number: 2x + x - 12 difference between -28 and the number: -28 - x those two are equal: 2x + x - 12 = -28 - x

OpenStudy (whpalmer4):

have to read what is there, not what you expect to be there :-)

OpenStudy (jhannybean):

Oh.. -28 MINUS x...yeah there we go... haha. ty!

OpenStudy (anonymous):

idk if this is right but I keep getting 40

OpenStudy (whpalmer4):

well, let's check it out. 2x 40 = 80 12 less 40 = -28 sum of those two is 80-28 = 52 difference between -28 and 40 = -28-40 = -68 -68 = 52 oops.

OpenStudy (whpalmer4):

Hint: the answer is a negative number..

OpenStudy (jhannybean):

2x + x -12 = -28 - x 3x - 12 = -28 - x 3x + x = -28 + 12 4x = -16 x = ?

OpenStudy (anonymous):

-4

OpenStudy (jhannybean):

thanks @whpalmer4 :)

OpenStudy (whpalmer4):

@kimbradunn let's check: twice the number: 2*-4 = -8 12 less than -4 = -8-12 = -16 sum of those two: -24 difference between -28 and -4 : -28 - (-4) = -28 + 4 = -24 good job!

OpenStudy (anonymous):

thanks everyone!!!